# Event shape dependence of symmetry plane correlations using the Gaussian estimator in Pb-Pb collisions at the LHC using AMPT

### v1

@@ IV Summary and Conclusion
In summary, we study centrality and transverse spherocity dependence of symmetry plane correlations in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using AMPT. For the first time, this paper presents a study of SPCs using AMPT and event shape dependence of SPCs. The Gaussian estimator of SPCs has been observed to be strongly sensitive to event shape, and their variation with spherocity can be used to study them in greater detail in different centrality classes of interest, as per the physics process being addressed. When SPCs are computed using isotropic events, they have been observed to have smaller magnitude and random variation about zero, showing no particular centrality dependence, except for $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}$, where, it can be distinctly observed that the SPC obtained using isotropic events only takes on negative values, irrespective of the centrality. Most of the SPCs, except for those that are very close to zero, when computed for the low-$S_{0}$ events, are all larger than the $S_{0}$-int case. Thus, these will be particularly useful in examining if a certain SPC is actually zero due to the underlying geometry of the system or if it is zero simply due to it being computed from the chosen sample of events with all kinds of event shapes. Symmetry planes of even orders are likely to have greater correlation, as has been seen for the SPCs of $n=2,4,6$. All the SPCs, especially for the low-$S_{0}$ case, qualitatively mimic the centrality dependence of their corresponding PPCs. The PPCs do not depend strongly on spherocity, as they are not affected by the several processes that ultimately lead spherocity to take on the value that it does for an event. The qualitative estimation of PPCs by the SPCs will be an important tool to determine the relevant final state effect of nuclear structure at high-energy heavy-ion collisions, such as in 16O-16O collisions at the RHIC and LHC.
@@ Acknowledgment
S. T. acknowledges the KVPY SB scholarship. S. P. acknowledges the University Grants Commission (UGC), Government of India, for doctoral fellowship. R. S. sincerely acknowledges the DAE-DST, Government of India funding under the mega-science project – “Indian participation in the ALICE experiment at CERN” bearing Project No. SR/MF/PS-02/2021-IITI (E-37123). The authors gratefully acknowledge the computing facility provided by the Grid Computing Facility, VECC, Kolkata, India.

### v2

@@ IV Summary and Conclusion
In summary, we study the centrality and transverse spherocity dependence of symmetry plane correlations in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using AMPT. For the first time, this paper presents a study of transverse spherocity dependence of SPCs using Gaussian estimators within the framework of AMPT. The Gaussian estimator of SPCs has been observed to be strongly sensitive to event shape, and their variation with spherocity can be used to study them in greater detail in different centrality classes of interest, as per the physics process being addressed. When SPCs are computed using isotropic events, they have been observed to have smaller magnitude and random variation about zero, showing no particular centrality dependence, except for $\langle\cos[4(\psi_{4}-\psi_{2})]\rangle_{\rm GE}$, where, it can be distinctly observed that the SPC obtained using isotropic events only takes on negative values, irrespective of the centrality.
Most of the SPCs, except for those that are very close to zero, when computed for the low-$S_{0}$ events, are all larger than the $S_{0}$-int case. Thus, these will be particularly useful in examining if a certain SPC is actually zero due to the underlying geometry of the system or if it is zero simply due to it being computed from the chosen sample of events with all kinds of event shapes. Symmetry planes of even orders are likely to have greater correlation, as has been seen for the SPCs of $n=2,4,6$. All the SPCs, especially for the low-$S_{0}$ case, qualitatively mimic the centrality dependence of their corresponding PPCs. The PPCs do not depend strongly on spherocity, as they are not affected by the several processes that ultimately lead spherocity to take on the value that it does for an event. The qualitative estimation of PPCs by the SPCs will be an important tool to determine the relevant final state effect of nuclear structure at high-energy heavy-ion collisions, such as in 16O-16O collisions at the RHIC and LHC [36].
@@ Acknowledgment
S. T. acknowledges the KVPY SB scholarship. S. P. acknowledges the University Grants Commission (UGC), Government of India, for doctoral fellowship. R. S. sincerely acknowledges the DAE-DST, Government of India, funding under the mega-science project – “Indian participation in the ALICE experiment at CERN” bearing Project No. SR/MF/PS-02/2021-IITI (E-37123). The authors gratefully acknowledge the computing facility provided by the Grid Computing Facility, VECC, Kolkata, India.